Constants Part A Three point charges are arranged on a line. Charge q3 +5.00 nC and is at the origin. Charge q2 -2.00 nC and is at 3.50 cm. Charge qi is at a - 1.50 cm What is qi (magnitude and sign) if the net force on g3 is zero? You may want to review (Pages 690-695) For related problemsolving tips and strategies, you may want to view a Video Tutor Solution of Vector addition of electric forces on a line. A2o 91 nC SubmitPrev Previous Answers Request Answer
Answer #1

If force experince by q3 is zero, it means force due to q1 is equal to force due to q2

k q1 q3 / r1^2 = k q2 q3 / r2^2

q1/ 1.5^2 = 2 / 3.5^2

q1 = 0.37 nC

Comment in case any doubt please rate my answer ...